\small
 \begin{thebibliography}{9}


 \bibitem{BIM}
 K.A. Bronnikov, V.D. Ivashchuk and V.N. Melnikov
 {\it Grav.  Cosmol. } {\bf 3}, 203-212 (1997); gr-qc/9710054.

 \bibitem{IMJ}
 V.D Ivashchuk, V.N. Melnikov, {\it J. Math. Phys. } {\bf 39},
 2866-2889 (1998); hep-th/9708157.

 \bibitem{Br1}
 K.A. Bronnikov, {\it Grav. Cosmol.} {\bf 4}, No 1 (13), 49 (1998);
 hep-th/9710207.

\bibitem{IMJ2}
 V.D. Ivashchuk and V.N. Melnikov,  In
 Lecture Notes in Physics, Vol. 537, "Mathematical and Quantum
 Aspects of Relativity and Cosmology Proceedings of the Second
 Samos Meeting on Cosmology, Geometry and Relativity held at
 Pythagoreon, Samos, Greece, 1998, eds: S. Cotsakis, G.W. Gibbons.,
 Berlin, Springer, pp. 214-247, 2000; gr-qc/9901001.

 \bibitem{CIM}
 S. Cotsakis, V.D. Ivashchuk and V.N. Melnikov, {\it Grav. Cosmol.}
 {\bf 5}, No 1 (17), 52-57 (1999); gr-qc/9902148.

 \bibitem{IMp1}
 V.D. Ivashchuk and V.N. Melnikov,
 {\it Grav.  Cosmol.} {\bf 5}, No 4 (20), 313-318
 (1999); gr-qc/0002085.

 \bibitem{IMp2}
 V.D. Ivashchuk and V.N. Melnikov, {\it Grav.  Cosmol.} {\bf 6}, No 1
(21), 27-40 (2000); hep-th/9910041.

 \bibitem{IMp3}
 V.D. Ivashchuk and V.N. Melnikov, {\it Class. Quantum
 Grav.} {\bf 17}, 2073-2092 (2000); math-ph/0002048.

 \bibitem{GrIvKim1}
  M.A. Grebeniuk, V.D. Ivashchuk and S.-W. Kim,
  {\it J. Math. Phys.} {\bf 43}, 6016-6023 (2002);
  hep-th/0111219.

 \bibitem{GrIvMel2}
   M.A. Grebeniuk, V.D. Ivashchuk and V.N. Melnikov
   {\it Phys. Lett. }, {\bf B 543}, 98-106
   (2002);  hep-th/0208083.

 \bibitem{IMtop}
    V.D. Ivashchuk and V.N. Melnikov, {\it Class. Quantum Grav. }
    {\bf 18} R87-R152 (2001); hep-th/0110274.

 \bibitem{IMS1}
    V.D. Ivashchuk, V.N. Melnikov and A.B. Selivanov,
    {\it Grav. Cosmol.} {\bf 7} No. 4 (12),
    (2001); gr-qc/0205103.

 \bibitem{IMS2}
    V.D. Ivashchuk, V.N. Melnikov and A.B. Selivanov,
    {\it Grav. Cosmol.} {\bf 9}, No. 1-2 (33-34), 50-54 (2003);
    hep-th/0211247.

  \bibitem{DIM-02}
   H. Dehnen, V.D. Ivashchuk and V.N. Melnikov,
   {\it Grav. Cosmol.} {\bf 9}, No. 3 (35), 153-158 (2003);
   gr-qc/0211049.

   \bibitem{DI-03}
   H. Dehnen and V.D. Ivashchuk,
   {\it J. Math. Phys.} {\bf 45}, 4726-4736 (2004);
   gr-qc/0310043.

  \bibitem{CT}
    M. Cvetic and A. Tseytlin,
    {\it Nucl. Phys.} {\bf B 478}, 181-198 (1996); hep-th/9606033.

  \bibitem{AIV}
    I.Ya. Aref'eva, M.G. Ivanov and I.V. Volovich, {\it
    Phys. Lett. } {\bf B 406}, 44-48 (1997); hep-th/9702079.

  \bibitem{Oh}
  N. Ohta, {\it Phys. Lett. } {\bf B 403}, 218-224 (1997);  hep-th/9702164.

  \bibitem{IK}
  V.D. Ivashchuk and S.-W. Kim, {\it J. Math. Phys.}, {\bf 41} (1)
   444-460 (2000); hep-th/9907019.

  \bibitem{IM5}
    V.D. Ivashchuk and V.N. Melnikov, {\it Int. J. Mod. Phys. }
    {\bf D 3}, No. 4, 795-811 (1994); gr-qc/ 9403063.

 \bibitem{GIM}
    V.R. Gavrilov, V.D. Ivashchuk, V.N. Melnikov,
   {\it J. Math. Phys } {\bf 36}, 5829--5847 (1995).

 \bibitem{IMZ}
    V.D. Ivashchuk, V.N.  Melnikov and A.I. Zhuk,
    {\it Nuovo Cim. } {\bf B 104}, No. 5, 575-581  (1989).

 \bibitem{York}
    J.W. York, {\it Phys. Rev. } {\bf D 31}, 775 (1985).

 \bibitem{FS}
 J. Fuchs and C. Schweigert, Symmetries, Lie algebras and
 Representations. A graduate course for physicists, Cambridge
 University Press, Cambridge, 1997.

 \bibitem{LMMP}
 H. L\"u, J. Maharana, S. Mukherji  and C.N. Pope,
  {\it Phys. Rev.} {\bf D 57 } 2219-2229 (1997); hep-th/9707182.

 \bibitem{I-98}
  V.D. Ivashchuk,
  {\it Phys. Lett.}, {\bf B 434}, No 1, 28-35  (1998);
   hep-th/9704113.

 \bibitem{I-10}
 V.D. Ivashchuk, in preparation.

 \bibitem{Mald}
 J.M. Maldacena, {\it Adv. in Theor.  Math. Phys.} {\bf
2}: 231–252 (1998); hep-th/9711200.

 \bibitem{GubKlPol}
S.S. Gubser, I.R. Klebanov and A.M. Polyakov,  {\it Phys. Lett.}
{\bf B 428}, 105–114 (1998); hep-th/9802109.

 \bibitem{Witten}
  E. Witten (1998), {\it Adv. in Theor.  Math.
Phys.} {\bf 2}, 253–291 (1998); hep-th/9802150.


 \end{thebibliography}

\documentclass[fleqn]{article}
\usepackage{amsfonts}

\newcommand{\bls}[1]{\renewcommand{\baselinestretch}{#1}}

\catcode`\@=11 \@addtoreset{equation}{section}\catcode`\@=12
\def\beq#1{\begin{equation}\label{#1}}
\def\eeq{\end{equation}}

 \def\bearr{\begin{eqnarray} \lal}
 \newcommand{\bear}[1]{\begin{eqnarray}\label{#1}}
 \def\ear{\end{eqnarray}}

 \def\lal{&&\nqq {}}
 
  \def\yy{\\[5pt] {}}
 \def\nqq{\hspace*{-2em}}
 \def\dys{\displaystyle}
 \def\mst{\mathstrut}

\def\e{{\,\rm e}}

 \newcommand{\fnm}{\footnotemark}
 \newcommand{\fnt}{\footnotetext}

 \newcommand{\R}{ {\mathbb R} }

 \bls{0.957}

\begin{document}

  \begin{center}

  \large \bf On analogues of black brane solutions  in the
        model  with multicomponent anisotropic fluid
  \end{center}

 \vspace{0.3truecm}

 \begin{center}

 \normalsize\bf V. D. Ivashchuk\fnm[1]\fnt[1]{e-mail:
  ivashchuk@mail.ru},

\vspace{0.3truecm}

 \it Center for Gravitation and Fundamental Metrology,
 VNIIMS, 46 Ozyornaya ul., Moscow 119361, Russia

 \it Institute of Gravitation and Cosmology,
 Peoples' Friendship University of Russia,
 6 Miklukho-Maklaya ul., Moscow 117198, Russia

\end{center}

\begin{abstract}

 A family of spherically symmetric solutions with horizon in the
model with $m$-component anisotropic fluid  is presented. The
metrics  are defined on a manifold that contains a product of
 $n-1$ Ricci-flat ``internal'' spaces. The equation of state for
any $s$-th component is defined by a vector $U^s$ belonging to
 $\R^{n + 1}$. The solutions  are governed by   moduli
functions $H_s$ obeying non-linear differential  equations with
certain boundary conditions imposed.  A simulation of black brane
solutions in the model with antisymmetric forms is considered. An
example of solution imitating  $M_2-M_5$ configuration (in $D =11$
supergravity) corresponding to Lie algebra  $A_2$  is presented.

\end{abstract}



\section{Introduction}

In this paper we continue our investigations of
spherically-symmetric solutions with horizon (e.g., black brane
ones) defined on product manifolds containing several Ricci-flat
factor-spaces (with diverse signatures and dimensions). These
solutions appear either in models with antisymmetric forms and
scalar fields \cite{BIM}-\cite{IMtop} or in models with
(multi-component) anisotropic fluid \cite{IMS1}-\cite{DI-03}. For
black brane solutions with $1$-dimensional factor-spaces (of
Euclidean signatures) see \cite{CT,AIV,Oh} and references therein.

These  and  more general brane cosmological and spherically
symmetric solutions were obtained by reduction of the field
equations to the Lagrange equations corresponding to Toda-like
systems \cite{IMJ,IK}. An analogous reduction for models with
multicomponent anisotropic fluids was performed earlier in
\cite{IM5,GIM}. For cosmological-type models with antisymmetric
forms without scalar fields any brane  is equivalent to an
anisotropic fluid with the equations of state:

\beq{0}
  \hat{p}_i =  - {\hat \rho}  \qquad
  {\rm or} \qquad {\hat p}_i = {\hat \rho},
 \eeq

when the manifold $M_i$ belongs or does not belong to the brane
worldvolume, respectively (here ${\hat p}_i$ is the effective
pressure in  $M_i$  and ${\hat \rho}$ is the effective density).

In this paper we  present spherically-symmetric solutions with
horizon (e.g the analogues of intersecting black brane solutions)
in a model with multi-component anisotropic fluid (MCAF), when
certain relations on fluid parameters are imposed. The solutions
are governed by  a set of moduli functions $H_s$ obeying
non-linear differential master equations with certain boundary
conditions imposed. These  master equations  are equaivalent to
Toda-like equations and depend upon the non-degenerate ($m \times
 m$)  matrix  $A$.  It was conjectured earlier that the functions $H_s$
 should be polynomials when $A$ is a Cartan matrix for some semisimple
 finite-dimensional Lie algebra (of rank $m$) \cite{IMp1}.  This conjecture
 was verified  for Lie algebras: $A_m$, $C_{m+1}$, $m \geq
 1$  \cite{IMp2,IMp3}.  A special case
 of black hole solutions with MCAF corresponding to semisimple Lie
 algebra $A_1 \oplus ... \oplus A_1$  was considered earlier in
\cite{IMS2} (for  $m = 1$ see \cite{IMS1}).

 The paper is organized as follows. In Section 2 the model is
 formulated. In Section 3  spherically-symmetric MCAF solutions
 with horizon corresponding to black brane type solutions,
 are presented.  In Section 4 a polynomial structure of
 moduli functions $H_s$ for semisimple finite-dimensional
 Lie algebras is discussed.
 In Section 5 a simulation of intersecting black brane solutions is
considered and an analogue  of $M2 - M5$ dyonic solution is
presented.

\section{The model}

In this paper we deal with a family of spherically symmetric
solutions to Einstein equations with an anisotropic matter source
 \beq{1.1}
        R^M_N - \frac{1}{2}\delta^M_N R = k^2 T^M_N,
 \eeq
defined on the manifold

\beq{1.2}
\begin{array}{l}
M = {\R}_{*}\times (M_{0}=S^{d_0}) \times (M_1 = {\R})
    \times \ldots \times M_n,\\ \qquad ^{\rm radial
    \phantom{p}}_{\rm variable}\quad^{\rm spherical}_{\rm variables}
            \quad\qquad^{\rm time}
\end{array}
\eeq
with the block-diagonal metrics

\beq{1.2a}
    ds^2= \e^{2\gamma (u)} du^{2}+\sum^{n}_{i=0}
            \e^{2 \beta^i(u)} h^{[i]}_{m_i n_i }dy^{m_i}dy^{n_i }.
\eeq

Here $\R_{*} \subseteq \R$ is an open interval. The manifold $M_i$
with the metric $h^{[i]}$, $i=1,2,\ldots,n$, is a Ricci-flat space
of dimension $d_{i}$:

 \beq{1.3}
    R_{m_{i}n_{i}}[h^{[i]}]=0,
 \eeq

and $h^{[0]}$ is the standard metric on the unit sphere $S^{d_0}$,
so that

 \beq{1.4}
    R_{m_{0}n_{0}}[h^{[0]}]=(d_0-1)h^{[0]}_{m_{0}n_{0}};
 \eeq

 $u$ is a radial variable, $\kappa^2$ is the gravitational constant,
 $d_1 = 1$ and $h^{[1]} = -dt \otimes dt$.

The energy-momentum tensor is adopted in the following form for
each component of the fluid:

\beq{1.5}
 ({T^{s}}^{M}_{N})= {\rm diag} (-{\hat{\rho}^{s}},
 {\hat p}_{0}^{s}
    \delta^{m_{0}}_{k_{0}}, {\hat p}_{1}^{s}
    \delta^{m_{1}}_{k_{1}},\ldots , {\hat p}_n^{s}
    \delta^{m_{n}}_{k_{n}}),
\eeq

where $\hat{\rho}^{s}$ and $\hat p_{i}^{s}$ are the effective
density and pressures respectively, depending on the radial
variable $u$.

We assume that the following ``conservation laws''
 \beq{5.0}
 \nabla_{M}T^{(s)M}_{\ N}=0 \eeq
 are valid for all components.

We also impose the following equations of state
\beq{1.7}
    {\hat p}_i^{s}=\left(1-\frac{2U_i^{s}}{d_i}\right){\hat{\rho}^{s}},
\eeq

where $U_i^{s}$ are constants, $i= 0,1,\ldots,n$.

The physical density and pressures are related to the effective
ones (with ``hats'') by the formulae
  \beq{1.7a}
        \rho^{s} = - {\hat p}_1^{s}, \quad p_u^{s} = - \hat{\rho}^{s},
        \quad p_i^{s} = \hat{p}_i^{s} \quad (i \neq 1).
   \eeq

In what follows we put $\kappa =1$ for simplicity.

\section{Spherically symmetric solutions with horizon}

We will make the following assumptions:

\beq{2.1}
\begin{array}{l}
    1^{o}.\quad U^{s}_0 = 0 \
    \Leftrightarrow \ \hat p^{s}_0 = \hat\rho^{s} ,\yy
    2^o.\quad U^{s}_1 = 1 \
    \Leftrightarrow \  \hat p^{s}_1 = -\hat\rho^{s} ,\yy
    3^o.\quad (U^{s},U^{s})  = U^{s}_i G^{ij}U^{s}_j > 0,
    \yy
    4^o.  \quad 2(U^{s},U^{l})/ (U^{l},U^{l}) = A_{sl},
\end{array}
\eeq

 where $A = (A_{sl})$ is non-degenerate matrix,

\beq{2.2a}
    G^{ij}=\frac{\delta^{ij}}{d_i} + \frac{1}{2-D},
\eeq are components of the matrix inverse to the matrix of the
minisuperspace metric \cite{IMZ}

 \beq{2.2}
    (G_{ij}) = (d_i \delta_{ij} - d_i d_j),
 \eeq

 $i,j = 0,1,...,n$ and $D=1 + \sum\limits_{i=0}^n {d_i}$ is the
total dimension.

 The conditions $1^o$ and $2^o$ in brane terms mean that brane
``lives'' in the time manifold $M_1$ and does not ``live'' in
 $M_{0}$. Due to  assumptions $1^o$ and $2^o$ and the equations of
state (\ref{1.7}),  the energy-momentum tensor (\ref{1.5}) reads
as follows:

\beq{2.1a}
    ({T^{(s)}}^{M}_{N})= {\rm diag} (-\rho^{s},\
    \rho^{s} \delta^{m_0}_{k_0},\
    - \rho^{s}, \ p_2^{s}\delta^{m_2}_{k_2},\ \ldots ,
    p_n^{s}\delta^{m_n}_{k_n}).
\eeq

Under the conditions (\ref{1.7}) and (\ref{2.1}) we have obtained
the following black-brane-like solutions to the Hilbert-Einstein
equations (\ref{1.1}):

\bear{12} \lal
    ds^{2} = J_{0}\left( \dys\frac{\mst
    dR^{2}}{1-{2\mu}/{R^{d}}} + R^{2} d \Omega^2_{d_0} \right) -
        J_1\left(1-\frac{2\mu}{R^{d}}\right)dt^{2}
    + \sum_{i=2}^{n} J_{i} h^{[i ]}_{m_{i}n_{i}} dy^{m_{i}}dy^{n_{i}},
 \\  \lal \label{13}
    \rho^{s}= - \frac{A_s}{J_0 R^{2d_0}} \prod_{l=1}^m H_l^{-A_{sl}} ,
    \ear
 which may be derived by analogy with the black brane solutions
 \cite{IMp2,IMp3}. Here
  $d=d_0-1$,
  \beq{2.sp}
    d \Omega_{d_0}^2= h^{[0]}_{m_{0}n_{0}} dy^{m_{0}}dy^{n_{0}}
  \eeq
is the $d_0$-dimensional spherical element (corresponding to the
metric on $S^{d_0}$),

\beq{2.3}
    J_{i} = \prod_{s =1}^m H_s^{-2 h_s U^{s i }},
\eeq
 $i = 0,1,...,n$,   $\mu >0$ is integration constant and
 \bearr\label{2.4}
    U^{s i} = G^{ij}U^{s}_{j}  = \frac{U^{s}_i}{d_i} + \frac{1}{2-D}
                \sum_{j=0}^{n}U^{s}_j,
  \\ \lal \label{2.4a}
        h_s = K_s^{-1}, \quad  K_s = ({U^{s}},{U^{s}}).
\ear
 It follows from  $1^{o}$ and (\ref{2.4}) that
   \beq{2.4b}
    U^{s 0} = \frac{1}{2-D} \sum_{j=1}^{n}U^{s}_j .
  \eeq

 Functions $H_s > 0$ obey the equations

 \beq{2.2.1}
  R^{d_0} \frac{d}{dR} \left[ \left(1 -
  \frac{2\mu}{R^{d}}\right) \frac{ R^{d_0} }{H_s} \frac{d
  H_s}{dR} \right] =  B_s \prod_{l = 1}^{m}  H_{l}^{- A_{s l}},
 \eeq
 with $B_s = 2 K_s A_s$
 and the boundary conditions imposed:
  \beq{2.2.1a}
   H_s  \to H_{s0} \neq 0, \quad
    {\rm for} \quad R^{d} \to  2\mu,
   \eeq
  and
  \beq{2.2.1b}
  H_s (R = +\infty) = 1,
  \eeq
  $s = 1,..., m$.

  Here we also impose  the following (additional) condition on the solutions
   \beq{2.2.1c}
     H_s(R) > 0  \ {\rm is \ smooth \ in} \ (R_{\epsilon}, +
     \infty),
   \eeq
  $s = 1,..., m$, where $R_{\epsilon}= (2 \mu)^{1/d} e^{- \epsilon}$,
  $\epsilon > 0$.  Then  the metric (\ref{12}) has a regular horizon
  at $R^{d} =   2 \mu$ and has an asymptotically flat
  $(2 + d_0)$-dimensional section.

   Due to to (\ref{2.1}) and (\ref{2.4}) the metric reads
   \beq{12a}
    ds^{2} = J_{0}\left[ \dys\frac{\mst
    dR^{2}}{1-{2\mu}/{R^{d}}} + R^{2} d \Omega^2_{d_0}  -
       \left(\prod_{s =1}^m H_s^{-2 h_s } \right)
       \left(1-\frac{2\mu}{R^{d}}\right)dt^{2}
    + \sum_{i=2}^{n} Y_{i} h^{[i ]}_{m_{i}n_{i}} dy^{m_{i}}dy^{n_{i}} \right],
    \eeq
  where
  \beq{2.3a}
  Y_i = \prod_{s =1}^m H_s^{-2 h_s U^{s}_i/d_i }.
  \eeq

  The solution (\ref{12}), (\ref{13}) may be verified just  by
  a straightforward substitution into equations of motion.
  A detailed derivation of this solution will be given in
  a separate paper \cite{I-10}. A special
  orthogonal case when $(U^s,U^{l})= 0$, for  $s \neq l$,
  was considered earlier in \cite{IMS2} (for $m =1$ see \cite{IMS1})
  More general solutions in orthogonal case (with more general condition
  instead of $2^o$) were obtained in  \cite{DI-03}
  (for $m =1$  see \cite{DIM-02}.)

 \section{Polynomial structure of $H_s$ for  Lie algebras}

Now we deal  with solutions to second order non-linear
differential equations  (\ref{2.2.1}) that may be rewritten as
follows

\beq{5.3.1}
 \frac{d}{dz} \left( \frac{(1 - 2\mu z)}{H_s}
 \frac{d}{dz} H_s \right) = \bar B_s
 \prod_{l =1}^{m}  H_{l}^{- A_{s l}}, \eeq

 where $H_s(z) > 0$,
  $z = R^{-d} \in (0, (2\mu)^{-1})$ ($\mu > 0$) and $\bar B_s =
 B_s/ d^2 \neq 0$. Eqs. (\ref{2.2.1a}) and  (\ref{2.2.1b})
 read

 \bear{5.3.2a}
  H_{s}((2\mu)^{-1} -0) = H_{s0} \in (0, + \infty), \\
 \label{5.3.2b} H_{s}(+ 0) = 1, \ear
  $s = 1,..., m$.

 The condition  (\ref{2.2.1c}) reads as follows
   \beq{5.3.2c}
     H_s(z) > 0 \ {\rm is \ smooth \ in} \ (0, z_{\epsilon}),
   \eeq
   $s = 1,..., m$, where $z_{\epsilon}= (2 \mu)^{-1} e^{ \epsilon d}$,
   $\epsilon > 0$.

   It was conjectured in \cite{IMp1}
 that  equations (\ref{5.3.1})-(\ref{5.3.2b})
 have  polynomial solutions  when  $(A_{s s'})$ is a  Cartan matrix for
 some  semisimple finite-dimensional Lie algebra $\cal G$ of rank
 $m$.  In this case we get

 \beq{5.3.12}
 H_{s}(z) = 1 + \sum_{k = 1}^{n_s} P_s^{(k)} z^k, \eeq

 where $P_s^{(k)}$ are constants, $k = 1,\ldots, n_s$;
 $P_s^{(n_s)} \neq 0$, and

 \beq{5.2.20}
  n_s = b_s \equiv
            2 \sum_{l = 1}^m  A^{s l}
  \eeq
 $s = 1,..., m$, are the components of twice the  dual Weyl
 vector in the basis of simple  co-roots \cite{FS}.
 Here $(A^{sl}) = (A_{sl}^{-1})$.

 This conjecture  was verified for ${\bf A_m}$ and ${\bf C_{m+1}}$
 series of Lie algebras in \cite{IMp2,IMp3}. In the extremal case ($\mu
 = + 0$) an analogue of this conjecture was suggested
(implicitly) in \cite{LMMP}.

 {\bf Remark.} We note that the substitution of (\ref{1.5}),
 (\ref{1.7}), (\ref{12}),  (\ref{13}) into Hilbert-Einstein
 equations (\ref{1.1}) gives an extra equation
 \beq{5.E}
  E_{T} = \frac{d^2}{4} \sum_{s,l =1}^m h_s A_{sl}
 \left[F \frac{d}{dz} \ln H_s + \mu b_s \right]
 \left[F \frac{d}{dz} \ln H_l + \mu b_l \right]
  +  \sum_{s =1}^m A_s F \prod_{l = 1}^{m}  H_{l}^{- A_{s l}} =
  \frac{1}{2} \sum_{s =1}^m h_s b_s (\mu d)^2,
 \eeq
 where $F = 1 - 2\mu z$. $E_{T}$ is an integral of motion
 for the set of equations (\ref{5.3.1}).
 The constraint (\ref{5.E})  is satisfied identically
 due to (\ref{5.3.2a}), (\ref{5.3.2c}) (one can check this by putting
 $2\mu z = 1$).

{\bf  ${\bf A_1} \oplus \ldots \oplus {\bf A_1}$ -case.}

The simplest example occurs in the orthogonal case : $(U^s,U^{l})=
0$, for  $s \neq l$ \cite{BIM,IMJ} (see also \cite{CT,AIV,Oh} and
refs. therein). In this case $(A_{s l}) = {\rm diag}(2,\ldots,2)$
is a Cartan matrix for the semisimple Lie algebra ${\bf A_1}
\oplus \ldots \oplus {\bf A_1}$ and

 \beq{5.3.5}
 H_{s}(z) = 1 + P_s z, \eeq
 with $P_s \neq 0$,  satisfying

 \beq{5.3.5a}
 P_s(P_s + 2\mu) = -\bar B_s = - 2 K_s A_s/d^2,
 \eeq
 $s = 1,..., m$. When all $A_s < 0$ (or, equivalently, $\rho^{s} > 0$) there
 exists a unique set of numbers  $P_s > 0$ obeying (\ref{5.3.5a}).

{\bf  $A_2$-case.}

 For the Lie algebra $\cal G$ coinciding with  ${\bf A_2} = sl(3)$
  we get $n_1 = n_2 =2$ and

 \beq{5.4.1} H_{s} = 1 + P_s z + P_s^{(2)} z^{2},
 \eeq

 where $P_s=
 P_s^{(1)}$ and $P_s^{(2)} \neq 0$ are constants, $s = 1,2$.

 It was found in \cite{IMp1} that for $P_1 +P_2 + 4\mu \neq 0$
 (e.g. when all $P_s >0 $) the following relations take place

 \bear{5.4.5}
 P_s^{(2)} = \frac{ P_s P_{s +1} (P_s + 2 \mu )}{2
(P_1 +P_2 + 4\mu)}, \qquad \bar B_s = - \frac{ P_s (P_s + 2 \mu
 )(P_s + 4 \mu )}{P_1 +P_2 + 4\mu}, \ear
 $s = 1,2$.

Here we denote $s+ 1 = 2, 1$ for $s = 1,2$, respectively.

 {\bf Other solutions.}

  At the moment the ``master'' equations were integrated (using
  Maple) in    \cite{GrIvKim1,GrIvMel2} for Lie
  algebras ${\bf C_2}$ and ${\bf A_3}$, respectively.
  (For ${\bf D_4}$-polynomials in the extremal case $\mu \to +0$ see \cite{LMMP}.)

  Special solutions $H_{s}(z) = (1 + P_s z)^{b_s}$
   with $b_s$ from (\ref{5.2.20})
  appeared earlier in  \cite{Br1,IMJ2,CIM} in a context of
  so-called block-orthogonal configurations.

  {\bf Extremal case.}
  For $\mu \to +0$  the conditions
  (\ref{5.3.2a}), (\ref{5.3.2c}) should be omitted
  but we should impose the relation
   \beq{5.Ee}
    E_{T} = \frac{d^2}{4} \sum_{s,l =1}^m h_s A_{sl}
    \frac{d}{dz} \ln H_s  \frac{d}{dz} \ln H_l
    +  \sum_{s =1}^m A_s F \prod_{l = 1}^{m}  H_{l}^{- A_{s l}} =
    0,
   \eeq
   following from  (\ref{5.E}). The functions $H_s(z) > 0$
   obeying (\ref{5.3.1}) (with $\mu \to +0$) are smooth on
   $(0,+ \infty)$.    For certain relations on
   $U^s$-vectors imposed the solution under consideration has a
   horizon for  $R \to +0$ ($z \to + \infty$), e.g. it may describe
   analogues  of extremal black brane solutions \cite{IMtop}.

   When the boundary condition (\ref{5.3.2b}) is omitted we get
  a special solution with

   \beq{5.nh}
          H_s(z) = C_s z^{b_s},
   \eeq
    and
    \beq{5.C}
          C_s = \prod_{l = 1}^{m} (-b_l/\bar B_l)^{-A^{sl}},
   \eeq
   where       $b_s/\bar B_s < 0$,
       $s = 1,..., m$. The metric (\ref{12}) with
       $H_s$ from (\ref{5.nh}) has no an asymptotically flat
       $(2 + d_0)$-dimensional section.

       It should be noted that  the  solutions obeying
        $H_{s}(+ 0) = 1$ and  $b_s/\bar B_s < 0$ have an asymptotical
       behaviour $H_s(z) \sim C_s z^{b_s}$ for
       $z \to + \infty$ (e.g. in the near-horizon  limit $R \to +0$
       of  extremal black-brane-type solutions).

  \section{Examples}

 \subsection{Analogues of intersecting black brane solutions}

The solution from the previous section for MCAF allows one to
simulate the intersecting black brane solutions \cite{IMtop} in
the model with antisymmetric forms without scalar fields. In this
case the parameters $U^{s}_i$ and pressures have  the following
form:

 \beq{3.1a}
 \begin{array}{ccccccr}
 U^{s}_i & = & d_i , & p_i^{s} & = & -\rho^{s} , &  i\in I_{s};\\
    &  & 0 , &  & & \rho^{s} ,  &  i \notin I_{s}.
 \end{array}
 \eeq

Here
 $I_{s} = \{ i_1^s, \ldots,  i_{k_s}^s \} \in \{1, \ldots n \}$
is the index set \cite{IMtop} corresponding to brane submanifold
 $M_{i_1^s}  \times \ldots \times M_{i_{k_s}^s}$.

The relation  $4^o$  (\ref{2.1}) leads us to  the following
dimensions of intersections of brane submanifolds
(``worldvolumes'')  \cite{IMJ,IMtop}:

  \beq{3.1b}
  d(I_s \cap I_l)=\frac{d(I_s)d(I_l)}{D-2} +
      \frac{1}{2}K_l A_{sl}, \eeq

 $s \neq l$; $s,l = 1,\ldots ,m$.
 Here $d(I_s)$ and $d(I_l)$ are dimensions of brane worldvolumes.

  \subsection{$M_2-M_5$-analogue   for Lie algebra $A_2$}

 In \cite{IMS2} examples of MCAF-analogues of
  $M2 \cap M5$, $M2 \cap M2$, $M5 \cap M5$ black brane solutions in
  $D = 11$ supergravity, with the standard (orthogonal) intersection
  rules  were considered.

 Now we consider a solution with 2-component anisotropic
 fluid that simulates $M_2-M_5$ dyonic configuration
  in $D = 11$ supergravity \cite{IMp1}, corresponding to Lie
 algebra ${\bf A_2}$.

 The solution is  defined on the manifold
\beq{5.4.8} M =  (2\mu, +\infty
  )  \times (M_0 = S^{2})  \times (M_1 = \R) \times M_{2} \times
 M_{3}, \eeq

 where ${\dim } M_2 =  2$ and ${\dim } M_3 =  5$.
 The $U^s$-vectors corresponding to fluid components
  obey  (\ref{3.1a}) with  $I_1 = \{ 1, 2 \}$ and $I_2 = \{ 1, 3 \}$.

The solution reads as following
 \begin{eqnarray}
  g =  H_1^{1/3} H_2^{2/3} \biggl\{ \frac{dR \otimes
  dR}{1 - 2\mu / R} + R^2   h[S^2]  -
  H_1^{-1} H_2^{-1} \left(1 - \frac{2\mu}{R} \right) dt\otimes dt
   + H_1^{-1} {h}^{[2]} + H_2^{-1} {h}^{[3]} \biggr\},
    \label{5.4.9} \\
     \rho^{1}= - \frac{A_1}{J_0 R^{4}}  H_1^{-2} H_2, \qquad
   \rho^{2}= - \frac{A_2}{J_0 R^{4}}  H_1 H_2^{-2},  \label{5.4.10}
 \end{eqnarray}

 where $J_0 = H_1^{1/3} H_2^{2/3}$; $h[S^2]$ is the canonical
 metric on 2-dimensional sphere $S^2$, $h^{[2]}$ and  $h^{[3]}$ are
 Ricci-flat metrics of Euclidean signatures defined on the manifolds
 $M_2$ and $M_3$, respectively; $\mu > 0$ and $H_s$
 are defined by (\ref{5.4.1}), where $z = R^{- 1}$ and parameters
 $P_s$, $P_s^{(2)}$, $\bar B_s  = B_s = 4 A_s$ ($s =1,2$)
 obey  (\ref{5.4.5}).

This  solution simulates ${\bf A_2}$-dyon from  \cite{IMp1}
consisting of an electric $M2$-brane with a worldvolume isomorphic
to $(M_1 = \R) \times M_{2}$ and a magnetic  $M5$-brane with a
 worldvolume isomorphic to $(M_1 = \R) \times M_{3}$. The branes
are intersecting on the time manifold $M_1 = \R$. Here $K_s =
(U^s,U^s)=2$,  for all $s =1,2$.

 For the ${\bf A_2}$-dyon from \cite{IMp1} we had $\bar B_s = B_s = - 2
 Q_s^2$, where $Q_s$ is the charge density parameter of $s$-th
 brane. Thus, for fixed $Q_s$ the fluid parameters should
 obey the relations $A_s = -\frac{1}{2} Q_s^2$ and hence
 $A_s$ are negative.

  Let us consider the extremal case $\mu \to +0$  of the solution
  (\ref{5.4.9})-(\ref{5.4.10}) with $A_1 < 0$ and $A_2 < 0$.
  The near-horizon limit ($R \to +0$) gives us an exact solution
  (see (\ref{5.nh}))
 \begin{eqnarray}
  g=  C_1^{1/3} C_2^{2/3} \biggl\{ h[AdS^2] +   h[S^2]
   + C_1^{-1} {h}^{[2]} +  C_2^{-1} {h}^{[3]} \biggr\},
 \label{5.4.9a} \\
  \rho^{1}= - A_1  C_1^{- 7/3} C_2^{1/3}, \qquad
   \rho^{2}= - A_2  C_1^{2/3} C_2^{-8/3},
   \label{5.4.10a}
 \end{eqnarray}
  where $C_1 = 2  |A_1|^{2/3} |A_2|^{1/3}$ and
  $C_2 = 2  |A_2|^{2/3} |A_1|^{1/3}$,
  $h[AdS^2] =   R^{-2}(dR \otimes  dR    -  R^4 d\bar t \otimes
  d \bar t)$ is the metric of (the half of) the anti-deSitter space $AdS^2$
   (here $\bar t =  C_1^{-1/2} C_2^{-1/2} t$).
   Thus, we have obtained a static configuration defined
   on (the half of) the product  space
   $AdS^2 \times S^2 \times M_2 \times M_3$. (For the
   solution with two branes and  $C_1 = C_2 =
   1$ see \cite{I-98}.)

  {\bf Analogues of  $M2$, $M5$ and $D3$  solutions.}
   Here we outline for completeness the analogues of non-marginal
   $M2$, $M5$ ($D = 11$) and $D3$ ($D = 10$)
   black brane solutions.
   The solutions are  defined on the product manifolds
   $M =  (2\mu, +\infty )  \times (M_0 = S^{d_0})  \times (M_1 = \R) \times
   M_{2}$,  where ${\dim } M_2 = d_2 = 2, 5, 3$ and $d_0 = 7, 4, 5 $
   for $M2$, $M5$, $D3$  branes, respectively.
   The vector $U = U^1$ has the components:
    $U_0 = 0$ and $U_i = 1$ for $i > 0$. The solutions read
    as follows
   \begin{eqnarray}
   g =  H^{r}  \biggl\{ \frac{dR \otimes dR}{1 - 2\mu R^{-d}}
   +  R^2   h[S^{d_0}]  -
    H^{-1}  \left(1 - 2\mu R^{-d} \right) dt\otimes dt
   + H^{-1} {h}^{[2]}  \biggr\},
    \label{5.4.11} \\
       \rho =  \rho^{1} = - A H^{-2 - r}  R^{ - 2 d_0}  \label{5.4.12},
 \end{eqnarray}
 where ${h}^{[2]}$ is a Ricci-flat metric on $M_2$,
 $h[S^{d_0}]$ is the canonical metric on $d_0$-dimensional sphere
 $S^{d_0}$,
  $H = 1 + P  R^{-d}$, $P(P + 2 \mu)= - 4 A/d^2 $ ($A < 0$, $P > 0$, $\mu >
 0$),  $d = d_0 -1$ and $r = 1/3, 2/3, 1/2$
  for $M2$, $M5$, $D3$  branes, respectively. In the
  extremal case  $\mu \to +0$  the near-horizon limit
  gives  an exact solution
  for the flat space $(M_2 = \R^{d_2},{h}^{[2]})$
  \begin{eqnarray}
  g=  P^{r} \biggl\{ h[S^{d_0}] + \frac{4}{(d -2)^2} h[AdS^{d_2 + 2}]
    \biggr\},
 \label{5.4.13} \\
  \rho = - A  P^{- 2 - r},  \qquad P^2 = - 4 A/d^2,
   \label{5.4.14}
 \end{eqnarray}
  where
  $h[AdS^{k + 2}] = du \otimes du +
  e^{2u}(- dy^0 \otimes dy^0 + \sum_{i =1}^k dy^i \otimes dy^i)$
  is the metric of (the part of) the anti-deSitter space
  $AdS^{k +2}$. Thus, we are led to a static configurations defined
   on (the parts of) the product  spaces
   $S^{d_0} \times AdS^{d_2 + 2}$, i.e.
   $S^{7} \times AdS^{4}$, $S^{4}\times AdS^{7}$
   and $S^{5} \times AdS^{5}$ for
    $M2$, $M5$ and $D3$  branes, respectively.
    It should be pointed out that the solutions with
    $AdS^{k}$ factor spaces may be of interest due
    to possible application in a context of
    $AdS/CFT$ approach \cite{Mald,GubKlPol,Witten}
    (or its modifications).

 \subsection{The Hawking temperature}

 The Hawking temperature of the black hole (\ref{12}) (see also
 (\ref{12a})) may be calculated using the relation from
 \cite{York}. It has the following form:

\beq{6.1}
 T_H=   \frac{d}{4 \pi (2 \mu)^{1/d}}
  \prod_{s = 1}^m H_{s0}^{- h_s}, \eeq

  where $H_{s0}$, $s =1,2$, are defined in
 (\ref{2.2.1a}).

For the dyonic solution from the previous subsection we get

  \beq{6.2}
 T_H =   \frac{1}{8 \pi  \mu}
   (H_{10}H_{20})^{- 1/2}, \eeq

 where $T_H$ is a function of fluid parameters $A_s < 0$,
 $s =1,2$.

\section{Conclusions}

Here we have presented a family of spherically symmetric solutions
with horizon in the model with multi-component anisotropic fluid
with the equations of state (\ref{1.7}) and the conditions
(\ref{2.1}) imposed.  The metric of any solution contains $(n -1)$
Ricci-flat ``internal'' space metrics.

As in  \cite{IMp1,IMp2,IMp3} the solutions are defined up to
solutions of non-linear differential  equations (equivalent to
Toda-like ones) with certain boundary conditions imposed. These
solutions may have a polynomial structure when the matrix $A$ from
 (\ref{2.1}) is coinciding with the Cartan matrix
 of some semi-simple finite-dimensional  Lie algebra.

For certain equations of state (with $p_i = \pm \rho$) the metric
of the solution  may coincide with the metric of intersecting
black branes (in a model with antisymmetric forms without
dilatons). Here we have considered an example of simulating of
$M2-M5$  black brane (dyonic) solution  in $D=11$ supergravity
with intersection rules corresponding to the Lie algebra $A_2$. We
have also outlined  the analogues of non-marginal $M2$, $M5$ and
$D3$  black brane solutions. In the extremal case $\mu \to +0$ the
near-horizon limits of all these solutions were found.

 An open problem is to generalize this formalism to the case when
 scalar fields are added into consideration.
 In a separate paper  we also plan to calculate the post-Newtonian
 parameters $\beta$ and $\gamma$ corresponding to the 4-dimensional
 section of the metric (for $d_0 =2$) and analyze the thermodynamic
 properties of the black-brane-like solutions in  the model with MCAF.

\begin{center}
 {\bf Acknowledgments}
 \end{center}

 This work was supported in part by
 grant NPK-MU (PFUR)
 and   Russian Foundation for Basic Research
 (Grant  Nr. 09-02-00677-a.).


 \end{document}